Situation 3: Melissa and Trevor sell candy bars to raise money for a class field trip. Trevor sells 1 more than five times as many as Melissa sells. Together they sell 49 candy bars.
Let m be the number of candy bars Melissa sells. Let t be the number of candy bars Trevor sells. Write a linear system to represent this situation.
49= t + m
T= 1+ 5m
If m is 5, t is 26. If m is 8 , t is 41 because Trevor sells 1 more than 5 times what Melissa sells.
I have done this much so far on my own. But I need help with finding the other linear equation with m. I have T but not m. Please show me how to set up M's equation.
thank you.

Answer by Boreal(15235)

t+m=49
t=5m+1
5m+1+m=49 by substitution
6m=48
m=8
t=41
You are correct all throughout. The linear system consists of the two equations above.
5m=t-1
m+(1/5)(t-1)